OpenStudy (anonymous):

points A and B are 100 apart and are of the same elevation as the foot a building. the angle of elevation of the top of the building from points A and B are 21 degrees and 32 degrees respectively. how far is A from the building in meters?

OpenStudy (anonymous):

A and B are 100 'meters?' apart?

OpenStudy (anonymous):

Well, here's a start: Let x = distance from A to building. Let y = distance from B to building. Let z = height of building. Set up two right triangles : tan(21º)=z/x, tan(32º)=z/y

OpenStudy (anonymous):

You'll need a third equation to fix x and y. Are you familiar with vectors? They may also come in handy.